OpenStudy (anonymous):

You have two block of clay in cube form and the edges are 10 cm. How many spheres with a radius of 5 cm can you make with that amount of clay?

OpenStudy (anonymous):

Find the total volume: 2 x 10 x 10 x 10 =

OpenStudy (anonymous):

It is 2000 cm3

OpenStudy (anonymous):

Volume of a sphere with a radius of 5 cm is =

OpenStudy (anonymous):

3 spheres

OpenStudy (anonymous):

that's the answer

OpenStudy (anonymous):

Meh lol correct

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

Np :P -formula- Volume of sphere w 5 cm radius=(4/3)(Pi)(r^3)=523.599 cc Volume 2 cubes clay w 10 cm sides=2(10)^3=2000cc You can make 3 spheres

OpenStudy (anonymous):

thanks again /.\

OpenStudy (anonymous):

np again :P